# SiReRAG: Indexing Similar and Related Information for Multihop Reasoning

### v1

@@ Abstract
Indexing is an important step towards strong performance in retrieval-augmented generation (RAG) systems. However, existing methods organize data based on either semantic similarity (similarity) or related information (relatedness), but do not cover both perspectives comprehensively. Our analysis reveals that modeling only one perspective results in insufficient knowledge synthesis, leading to suboptimal performance on complex tasks requiring multihop reasoning. In this paper, we propose SiReRAG, a novel RAG indexing approach that explicitly considers both similar and related information. On the similarity side, we follow existing work and explore some variances to construct a similarity tree based on recursive summarization. On the relatedness side, SiReRAG extracts propositions and entities from texts, groups propositions via shared entities, and generates recursive summaries to construct a relatedness tree. We index and flatten both similarity and relatedness trees into a unified retrieval pool. Our experiments demonstrate that SiReRAG consistently outperforms state-of-the-art indexing methods on three multihop datasets (MuSiQue, 2WikiMultiHopQA, and HotpotQA), with an average 1.9% improvement in F1 scores. As a reasonably efficient solution, SiReRAG enhances existing reranking methods significantly, with up to 7.8% improvement in average F1 scores.
@@ 1 Introduction
We show that SiReRAG is effective on a variety of multihop question answering (QA) datasets including MuSiQue (Trivedi et al., 2022), 2WikiMultiHopQA (Ho et al., 2020), and HotpotQA (Yang et al., 2018). We find that SiReRAG consistently outperforms the strongest RAG indexing methods, achieving an average F1 score improvement of at least 1.9%. We conduct an ablation study on several components in SiReRAG and observe that adding proposition aggregates to the similar information within a dataset yields the most improvement, which echoes our motivation. SiReRAG is also a reasonably efficient model, as it does not introduce many lengthy or redundant retrieval candidates.
@@ 4.2 Synthesizing information based on relatedness
We define a proposition as “a factual statement describing important information (preferably about some entities) from a paragraph”. We extract entities and propositions using the Distill-SynthKG pipeline (Choubey et al., 2024), adapting its SynthKG workflow. Specifically, we first rewrite chunks of documents taken from 10K BAAI/IndustryCorpus documents  [^2] https://huggingface.co/datasets/BAAI/IndustryCorpus to resolve entity references. Here we use Meta-Llama-3-70B-Instruct  [^3] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct (AI@Meta, 2024), and the LLM prompt of rewriting chunks is in Figure 6. Then, after we obtain entities from these rewritten chunks, we extract all relevant propositions and their associated entities (using the same LLM) as detailed in the LLM prompt in Figure 7. Finally, we consolidate all propositions and entities as the training data to fine-tune Mistral-7B-Instruct-v0.3  [^4] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 (Jiang et al., 2023). This smaller fine-tuned model is used to extract propositions and their associated entities from our multihop datasets.
@@ 4.3 Indexing similarity and relatedness trees
We propose to construct similarity and relatedness trees independently. This approach ensures that summary nodes in one tree do not access the clusters of the other, leading to a simpler design. There is another slightly complex design in which we allow summary nodes from one tree to access clusters from the other tree. This interaction may enables summary nodes in both trees to inform and enhance each other and improving their informativeness and consequently performance. However, this approach sacrifices the distinction between similarity and relatedness. Additionally, allowing cross-tree interaction leads to more nodes to cluster at each level as well as requires summarization based on a greater number of nodes per cluster, all of which increases the overall complexity of the system. We experimented with both settings and did not observe performance improvement as shown in Appendix A. Therefore, we opted for the simpler first implementation in our evaluation.
@@ 5.2 Baselines
We select RAPTOR, HippoRAG, and GraphRAG as state-of-the-art retrieval baselines. As discussed above, RAPTOR integrates knowledge based on similarity while the other two approaches focus on relatedness. Specifically, HippoRAG has both indexing and retrieval components, and we use ColBERTv2 (Santhanam et al., 2022) as the retriever of HippoRAG due to its strongest QA performance reported. Although GraphRAG has a different goal (global questions directed at an entire dataset) than ours, we include it to show its performance on multihop QA datasets. Since the queries in our datasets ask fine-grained details, we use the local search function of GraphRAG instead of its global search.
@@ 6.1 Overall results
As for GraphRAG, it considers relatedness solely, and it delivers the worst performance scores on our datasets. After a manual verification, we find that GraphRAG often provides “I don’t know” answers, suggesting that it prefers not to give a concrete answer. Since GraphRAG is designed to handle query-focused summarization of an entire corpus, it is not a competitive method in terms of accuracy for existing multihop QA tasks.
@@ Appendix E Performance on Single-Hop QA, MultiHop-RAG, and Ambiguous Questions
Appendix F LLM Prompts
The prompt we use to perform summarization on a cluster of nodes is “summarize the provided text, including as many key details as needed”. This prompt is the same as RAPTOR. In Section 3, we use ‘identify the high-level topic of this paragraph as concise as possible” to extract the topic of each passage. As mentioned in Section 4.1, the prompt used for identifying a two-level hierarchy for all chunks is shown in Figure 5. As mentioned in Section 4.2, the LLM prompt for rewriting chunks is shown in Figure 6, and the prompt for extracting propositions and entities is shown in Figure 7.
@@ Appendix F LLM Prompts
Figure 7: Prompt for extracting propositions and entities from a document.

### v2

@@ Abstract
Indexing is an important step towards strong performance in retrieval-augmented generation (RAG) systems. However, existing methods organize data based on either semantic similarity (similarity) or related information (relatedness), but do not cover both perspectives comprehensively. Our analysis reveals that modeling only one perspective results in insufficient knowledge synthesis, leading to suboptimal performance on complex tasks requiring multihop reasoning. In this paper, we propose SiReRAG, a novel RAG indexing approach that explicitly considers both similar and related information. On the similarity side, we follow existing work and explore some variances to construct a similarity tree based on recursive summarization. On the relatedness side, SiReRAG extracts propositions and entities from texts, groups propositions via shared entities, and generates recursive summaries to construct a relatedness tree. We index and flatten both similarity and relatedness trees into a unified retrieval pool. Our experiments demonstrate that SiReRAG consistently outperforms state-of-the-art indexing methods on three multihop datasets (MuSiQue, 2WikiMultiHopQA, and HotpotQA), with an average 1.9% improvement in F1 scores. As a reasonably efficient solution, SiReRAG enhances existing reranking methods significantly, with up to 7.8% improvement in average F1 scores. Our code is available at https://github.com/SalesforceAIResearch/SiReRAG.
@@ 1 Introduction
We show that SiReRAG is effective on a variety of multihop question answering (QA) datasets including MuSiQue (Trivedi et al., 2022), 2WikiMultiHopQA (Ho et al., 2020), and HotpotQA (Yang et al., 2018). We find SiReRAG consistently outperforms the strongest RAG indexing methods, achieving an average F1 improvement of at least 1.9%. We conduct an ablation study on several components in SiReRAG and observe that adding proposition aggregates to the similar information within a dataset yields the most improvement, which echoes our motivation. SiReRAG is also a reasonably efficient model, as it does not introduce many lengthy or redundant retrieval candidates.
@@ 4.2 Synthesizing information based on relatedness
We define a proposition as “a factual statement describing important information (preferably about some entities) from a paragraph”. We extract entities and propositions using the Distill-SynthKG pipeline (Choubey et al., 2024), adapting its SynthKG workflow. First, we rewrite chunks of 10K documents from the BAAI/IndustryCorpus [^2] https://huggingface.co/datasets/BAAI/IndustryCorpus to resolve entity references, using Meta-Llama-3-70B-Instruct [^3] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct (AI@Meta, 2024) with the rewriting prompt shown in Figure 6. Next, we prompt the same LLM to extract entities from these rewritten chunks (prompt is shown in Figure 7). After obtaining these entities, we again prompt the LLM to identify all relevant propositions and their associated entities (prompt is shown in Figure 8). We then consolidate the resulting propositions and entities to fine-tune Mistral-7B-Instruct-v0.3444https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 (Jiang et al., 2023). This smaller fine-tuned model is subsequently used to extract propositions and their associated entities from our multihop datasets.
@@ 4.3 Indexing similarity and relatedness trees
We propose to construct similarity and relatedness trees independently. This approach ensures that summary nodes in one tree do not access the clusters of the other, leading to a simpler design. There is another slightly complex design in which we allow summary nodes from one tree to access clusters from the other tree. This interaction may enable summary nodes in both trees to inform and enhance each other, improving their informativeness and consequently performance. However, this approach sacrifices the distinction between similarity and relatedness. Additionally, allowing cross-tree interaction leads to more nodes to cluster at each level as well as requires summarization based on a greater number of nodes per cluster, all of which increases the overall complexity of the system. We experimented with both settings and did not observe performance improvement as shown in Appendix A. Therefore, we opted for the simpler first implementation in our evaluation.
@@ 5.2 Baselines
We select RAPTOR, HippoRAG, and GraphRAG as state-of-the-art retrieval baselines. As discussed above, RAPTOR integrates knowledge based on similarity while the other two approaches focus on relatedness. Specifically, HippoRAG has both indexing and retrieval components, and we use ColBERTv2 (Santhanam et al., 2022) as the retriever of HippoRAG due to its strongest QA performance reported. Although GraphRAG has a different goal (global questions directed at an entire dataset) than ours, we include it to show its performance on multihop QA datasets. Since the queries in our datasets ask fine-grained details, we use the local search function of GraphRAG instead of its global search. Additional details of GraphRAG are specified in Appendix F.
@@ 6.1 Overall results
As for GraphRAG, it considers relatedness solely, and it delivers the worst performance scores on our datasets. After a manual verification, we find that GraphRAG often provides “I don’t know” answers, suggesting that it prefers not to give a concrete answer. Since GraphRAG is designed to handle query-focused summarization of an entire corpus, it is not the most competitive approach in terms of accuracy for existing multihop QA tasks.
@@ Appendix E Performance on Single-Hop QA, MultiHop-RAG, and Ambiguous Questions
Appendix F Additional GraphRAG Details
We adhere to GraphRAG documentation [^5] https://microsoft.github.io/graphrag/ to construct its indexing structure and handle QA. Following the recommendation, we use its “auto tuning” to generate domain adapted prompts for the creation of its knowledge graph. We use the default values for most hyperparameters, except that we set the response type to “a few words” and do not include covariates during indexing.
Appendix G LLM Prompts
The prompt we use to perform summarization on a cluster of nodes is “summarize the provided text, including as many key details as needed”. This prompt is the same as RAPTOR. In Section 3, we use ‘identify the high-level topic of this paragraph as concise as possible” to extract the topic of each passage. As mentioned in Section 4.1, the prompt used for identifying a two-level hierarchy for all chunks is shown in Figure 5. As mentioned in Section 4.2, the LLM prompt for rewriting chunks is shown in Figure 6, the prompt for extracting named entities from rewritten chunks is shown in Figure 7, and the prompt for extracting propositions is shown in Figure 8.
@@ Appendix G LLM Prompts
{mdframed}
Extract all named entities from the document. Also generate the type for each entity.
Instructions
- Generate only the most informative name for each named entity. Example: if John P., Parker, John Parker are coreferential, only generate John Parker.
- Use your best understanding best on the domain of paragraph to decide appropriate entity types.
- Respond using json format provided below.
    "n1":{"name": "entity_name", "type": "entity_type_label"},
    "n2":{},
Below is an example for reference.
Paragraph: Tucked into Eli Lilly’s year-end earnings report, the company revealed positive results from Synergy-NASH—its phase 2 study of tirzepatide in adults in nonalcoholic steatohepatitis (NASH), also known as metabolic dysfunction-associated steatohepatitis (MASH).
Output:
    "n1": {"name": "Eli Lilly", "type": "Organization"},
    "n2": {"name": "Synergy-NASH", "type": "Clinical Trial"},
    "n4": {"name": "tirzepatide", "type": "Drug"},
    "n5": {"name": "nonalcoholic steatohepatitis", "type": "Disease"},
    "n6": {"name": "metabolic dysfunction-associated steatohepatitis", "type": "Disease"},
    "n7": {"name": "year-end earnings report", "type": "Document"}
Figure 7: Prompt for extracting entities from a document (e.g., a rewritten chunk).
{mdframed}
Extract all facts from the document. For each fact, also generate all semantic triplets.
Instructions
- Consistently use the most informative name for each named entity in all facts and triplets.
- Avoid pronouns or ambiguous references in facts and triplets. Instead, directly include all relevant named entities in facts.
- Ensure that each semantic triplet contains head entity, predicate, and tail entity.
- Ensure that at least one (preferably both) entity in each semantic triplet is present in the given entities list.
- Respond using json format provided below:
    "f1":{
        "fact": "A factual statement describing important information (preferably about some entities) from the paragraph",
        "triplets: [["entity 1", "predicate", "entity 2"], ["entity 1", "predicate", "entity 3"]]
    "f2":{},
Below is an example for reference.
Paragraph: Locked in a heated battle with Novo Nordisk’s semaglutide franchise, Eli Lilly’s tirzepatide is beginning to come into its own—both with regards to sales and amid attempts to show the dual GIP/GLP-1 agonist can strike out beyond diabetes and obesity. As Mounjaro, tirzepatide won its first FDA nod in Type 2 diabetes back in May 2022. An obesity approval followed last November, with that formulation of tirzepatide adopting the commercial moniker Zepbound. In 2023’s fourth quarter, Mounjaro generated a whopping \$2.2 billion in sales, a nearly eight-fold increase over the \$279 million it pulled down during the same stretch in 2022. Year-to-date, the drug brought home around \$5.2 billion in revenues, Lilly said in an earnings release Tuesday. Zepbound, for its part, generated \$175.8 million during its first quarter on the market. Overall, Lilly reeled in around \$9.4 billion in fourth-quarter sales, growing 28% over the \$7.3 billion it made for the quarter in 2022.
Entities: Eli Lilly, Novo Nordisk, Tirzepatide, Semaglutide, GLP-1, GIP, FDA, Mounjaro, Zepbound
Output:
    "f1": {
        "fact": "Eli Lilly’s tirzepatide is competing with Novo Nordisk’s semaglutide franchise.",
        "triplets": [["Eli Lilly", "competing with", "Novo Nordisk"], ["Tirzepatide", "is competing with", "Semaglutide"]]
    "f2": {
        "fact": "Eli Lilly is trying to show tirzepatide, the dual GIP/GLP-1 agonist, can strike out beyond diabetes and obesity.",
        "triplets": [["Eli Lilly", "is trying to show", "Tirzepatide"], ["Tirzepatide", "is a", "dual GIP/GLP-1 agonist"],
                     ["Tirzepatide", "can treat beyond", "Diabetes"], ["Tirzepatide", "can treat beyond", "Obesity"]]
    "f3": {
        "fact": "Tirzepatide, under the brand name Mounjaro, received its first FDA approval for Type 2 diabetes in May 2022.",
        "triplets": [["Tirzepatide", "branded as", "Mounjaro"], ["Mounjaro", "won", "FDA approval"],
                     ["FDA approval", "for",  "Type 2 diabetes"], ["FDA approval", "was in", "May 2022"]]
    "f4": {
        "fact": "Tirzepatide, under the brand name Zepbound, received an obesity approval in November 2022.",
        "triplets": [["Tirzepatide", "was branded as", "Zepbound"], ["Zepbound", "received", "Obesity approval"],
                     ["Obesity approval", "was in", "November 2022"]]
    "f5": {
        "fact": "Mounjaro generated \$2.2 billion in sales in the fourth quarter of 2023, an eight-fold increase from the \$279 million
        during the same period in 2022.",
        "triplets": [["Mounjaro", "2023’s fourth quarter sales", "\$2.2 billion sales"],
        ["Mounjaro", "2022’s fourth quarter sales", "\$279 million"]]
    "f6": {
        "fact": "Mounjaro brought in around \$5.2 billion in revenues year-to-date in 2023, Lilly said in an earnings release Tuesday",
        "triplets": [["Mounjaro", "2023 sales year-to-date", "\$5.2 billion revenues"]]
    "f7": {
        "fact": "Zepbound generated \$175.8 million in sales in its first quarter on the market.",
        "triplets": [["Zepbound", "first quarter sales", "\$175.8 million"]]
    "f8": {
        "fact": "Eli Lilly’s fourth-quarter sales were around \$9.4 billion, a 28% increase over the \$7.3 billion during the same
        period in 2022.",
        "triplets": [["Eli Lilly", "2023 fourth-quarter sales", "\$9.4 billion,"],
        ["Eli Lilly", "2022 fourth-quarter sales", "\$7.3 billion,"]]
Figure 8: Prompt for extracting propositions (e.g., facts) and their corresponding entities from a document.